OpenStudy (anonymous):

d=vt+1/2at^2 solve for a

OpenStudy (across):

\[d=vt+\frac{1}{2}at^2,\]\[d-vt=\frac{1}{2}at^2,\]\[2(d-vt)=at^2,\]\[\frac{2}{t^2}(d-vt)=a.\]